OpenStudy (anonymous):

f-2x-5 how do i solve this show me how to do the steps

jhonyy9 (jhonyy9):

so i think that this wann being f(x)=2x-5 yes ?

jhonyy9 (jhonyy9):

so than make f(x)=0 than will get 2x-5=0 ok ?

jhonyy9 (jhonyy9):

so for the first step add to both sides 5 and will get 2x-5+5=0+5 so 2x=5 ---- can you continue it ?

jhonyy9 (jhonyy9):

ok so now divide both sides by 2 and will get 2x/2 = 5/2 x= 5/2

jhonyy9 (jhonyy9):

do you understand it sure ?

jhonyy9 (jhonyy9):

ok you dont like cooperat nothing good luck bye

OpenStudy (anonymous):

so sorry i wasnt looking

jhonyy9 (jhonyy9):

ok so do you understand it ?

OpenStudy (anonymous):

trying now to understand it hang on k

OpenStudy (anonymous):

yes thank you

jhonyy9 (jhonyy9):

ok good luck bye

OpenStudy (anonymous):

ty

jhonyy9 (jhonyy9):

welcom